OpenStudy (pphalke):

The formula m=1/3t can be used to determine how far M, in miles, you are from lightning when its thunder, takes t seconds to reach your ears. if it takes 3 sec for the sound of thunder to reach you after you have seen the lightning how far away is the storm

OpenStudy (jdoe0001):

m = 1/3 t pretty much, t = 3, how far is the storm, or what is "m"

OpenStudy (texaschic101):

m = 1/3 * 3 m = 1 mile

OpenStudy (texaschic101):

baically, u sub in 3 for t and solve for m

OpenStudy (texaschic101):

typo...basically